# A Study of stable wormhole solution with non-commutative geometry in the framework of linear $f(R,\mathcal{L}_{m},T)$ gravity

### v5

@@ III.1 Gaussian distribution
$\bm{\beta+\frac{\alpha}{2}<\frac{(4\pi\Theta)^{3/2}}{Mr_{0}^{2}}e^{\frac{r_{0}^{2}}{4\Theta}}-8\pi}.$ (31)
@@ III.2 Lorentzian distribution
$\bm{\beta+\frac{\alpha}{2}<\frac{\pi^{2}(\Theta+r_{0}^{2})^{2}}{\sqrt{\Theta}Mr_{0}^{2}}-8\pi,}$
@@
Figure 2: The figure illustrates the conditions of the shape function relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@
Figure 3: The figure illustrates the conditions of the shape function relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@ IV.1 Gaussian Distribution
$\bm{\beta+\frac{\alpha}{2}<-\frac{(4\pi\Theta)^{3/2}}{Mr_{0}^{2}}e^{\frac{r_{0}^{2}}{4\Theta}}-8\pi,}$ (45)
@@
Figure 4: The figure illustrates the profile of energy density and NEC relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@
Figure 5: The figure illustrates the profile of energy density and NEC relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@ IV.2 Lorentzian Distribution
$\bm{\beta+\frac{\alpha}{2}<-\frac{\pi^{2}(\Theta+r_{0}^{2})^{2}}{\sqrt{\Theta}Mr_{0}^{2}}-8\pi.}$
@@ V Gravitational Lensing Effects
the potential will have a maximum exactly at $r=r_{0}(l=0)$ [102]. Figures 6 and 8 demonstrate this behavior for both the Gaussian and Lorentzian mass distributions for different values of $r_{0}$ and $\beta$ and the usual choice of additional parameters. Thus, the throat itself must serve as an unstable photon sphere. This fact can be further investigated by considering the angle $\gamma$ by which the photon gets deflected:
@@
i.e., $\gamma$ can also be considered a function of the ratio $L/E$. As $r=r_{0}$ is the sole photon sphere of this system, the only singularity in Eq. (52) arises at $r_{tp}=r_{0}$ and for, $r_{tp}<r_{0}$, the function will diverge. The following sections will discuss the numerical results for different values of our parameters $\alpha$ and $\beta$.
Figure 6: The figure illustrates the profile of effective potential relative to the proper radial coordinate $l$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant
Figure 7: The figure illustrates the profile of light deflection angle $\gamma$ relative to the radial coordinate $r_{tp}$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, and $M=1$, are held constant.
@@ V.1 Gaussian Distribution
Figure 8: The figure illustrates the profile of effective potential relative to the proper radial coordinate $l$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
Figure 9: The figure illustrates the profile of light deflection angle relative to the radial coordinate $r_{tp}$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, and $M=1$, are held constant.
@@ VI Tolman-Oppenheimer-Volkoff Condition
Figure 10: The figure illustrates the changes in the anisotropic and hydrostatic forces relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@
Figure 11: The figure illustrates the changes in the anisotropic and hydrostatic forces relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,L_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@ VII Conclusion
$\bm{\beta+\frac{\alpha}{2}<\frac{(4\pi\Theta)^{3/2}}{Mr_{0}^{2}}e^{\frac{r_{0}^{2}}{4\Theta}}-8\pi}$ (63)
@@
$\bm{\beta+\frac{\alpha}{2}<\frac{\pi^{2}(\Theta+r_{0}^{2})^{2}}{\sqrt{\Theta}Mr_{0}^{2}}-8\pi}$ (64)

### v6

@@ III.1 Gaussian distribution
$\beta+\frac{\alpha}{2}<\frac{(4\pi\Theta)^{3/2}}{Mr_{0}^{2}}e^{\frac{r_{0}^{2}}{4\Theta}}-8\pi.$ (31)
@@ III.2 Lorentzian distribution
$\beta+\frac{\alpha}{2}<\frac{\pi^{2}(\Theta+r_{0}^{2})^{2}}{\sqrt{\Theta}Mr_{0}^{2}}-8\pi,$
@@
Figure 2: The figure illustrates the conditions of the shape function relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@
Figure 3: The figure illustrates the conditions of the shape function relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@ IV.1 Gaussian Distribution
$\beta+\frac{\alpha}{2}<-\frac{(4\pi\Theta)^{3/2}}{Mr_{0}^{2}}e^{\frac{r_{0}^{2}}{4\Theta}}-8\pi,$ (45)
@@
Figure 4: The figure illustrates the profile of energy density and NEC relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@
Figure 5: The figure illustrates the profile of energy density and NEC relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@ IV.2 Lorentzian Distribution
$\beta+\frac{\alpha}{2}<-\frac{\pi^{2}(\Theta+r_{0}^{2})^{2}}{\sqrt{\Theta}Mr_{0}^{2}}-8\pi.$
@@ V Gravitational Lensing Effects
the potential will have a maximum exactly at $r=r_{0}(l=0)$ [102]. Figures 6 and 8 demonstrate this behavior for both the Gaussian and Lorentzian mass distributions for different values of $r_{0}$ and $\beta$ and the usual choice of additional parameters. Thus, the throat itself must serve as an unstable photon sphere. This fact can be further investigated by considering the angle $\bm{\alpha}$ by which the photon gets deflected:
@@
i.e., $\bm{\alpha}$ can also be considered a function of the ratio $L/E$. As $r=r_{0}$ is the sole photon sphere of this system, the only singularity in Eq. (52) arises at $r_{tp}=r_{0}$ and for, $r_{tp}<r_{0}$, the function will diverge. The following sections will discuss the numerical results for different values of our parameters $\alpha$ and $\beta$.
Figure 6: The figure illustrates the profile of effective potential relative to the proper radial coordinate $l$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant
Figure 7: The figure illustrates the profile of light deflection angle ($\bm{\alpha}$) relative to the radial coordinate $r_{tp}$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, and $M=1$, are held constant.
@@ V.1 Gaussian Distribution
Figure 8: The figure illustrates the profile of effective potential relative to the proper radial coordinate $l$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
Figure 9: The figure illustrates the profile of light deflection angle ($\bm{\alpha}$) relative to the radial coordinate $r_{tp}$ for different values of $r_{0}$, $\alpha$, and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, and $M=1$, are held constant.
@@ VI Tolman-Oppenheimer-Volkoff Condition
Figure 10: The figure illustrates the changes in the anisotropic and hydrostatic forces relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Gaussian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@
Figure 11: The figure illustrates the changes in the anisotropic and hydrostatic forces relative to the radial coordinate $r$ for different values of $\alpha$ and $\beta$ [showing GR $\left(\alpha=0,\beta=0\right)$, $f(R,T)$ $\left(\alpha=0,\beta=1\right)$, and $f(R,\mathcal{L}_{m},T)$ $\left(\alpha=1,\beta\in\{{1,3\}}\right)$ cases] under the Lorentzian distribution. The other parameters, specifically $\Theta=2$, $M=1$, and $r_{0}=1$, are held constant.
@@ VII Conclusion
$\beta+\frac{\alpha}{2}<\frac{(4\pi\Theta)^{3/2}}{Mr_{0}^{2}}e^{\frac{r_{0}^{2}}{4\Theta}}-8\pi$ (63)
@@
$\beta+\frac{\alpha}{2}<\frac{\pi^{2}(\Theta+r_{0}^{2})^{2}}{\sqrt{\Theta}Mr_{0}^{2}}-8\pi$ (64)